OpenStudy (anonymous):

I just really don't understand how to solve this: g(u) = 3/(u-2), find and simplify [g(x+h) - g(x)]/h. the fraction is just throwing me off

OpenStudy (dumbcow):

ok so instead of u you plug in (u+h) \[g(u+h) = \frac{3}{(u+h)-2}\]

OpenStudy (anonymous):

g(x+h) = 3/(x+h-2) g(x) = 3/(x-2) so, g(x+h) - g(x) = -3h / (x+h-2)(x-2) so, (g(x+h) - g(x))/h = -3/(x+h-2)(x-2)

OpenStudy (dumbcow):

yeah like that ^^ what they did was combine the fractions with common denominator being (u+h-2)(u-2)

OpenStudy (dumbcow):

This is what the numerator would look like 3(u-2) - 3(u+h-2) = 3u-3u-6+6 -3h = -3h

OpenStudy (anonymous):

Thanks so much! You guys helped a ton